Question 70

The __________________ measures the rate of income earned on the amount invested by stockholders.

a.

Earnings per Share on Common Stock

b.

Rate Earned on Total Assets

c.

Rate Earned on Stockholders Equity

d.

Rate of net Sales to Assets

Question 86

Current Assets total $35,000 and Current Liabilities total $32,000. Of the Current Assets, Cash totals $13,000 and Accounts Receivable, $5,000. What is the Current Ratio? the Quick Ratio?

a.

Current Ratio=1.09

Quick Ratio=.56

b.

Current Ratio=1.09

Quick Ratio=.41

c.

Current Ratio=.51

Quick Ratio=.91

d.

Current Ratio=.91

Quick Ratio=.51
Answer #1

Q-70:-

The Rate Earned on Stockholders Equity measures the rate of income earned on the amount invested by stockholders.

Q-86:-

Current ratio =CA/CL=35000/32000 =1.09

Quick ratio =(13000+5000)/32000 =0.56